# Gravity as a topological gauge theory

### v1

@@ A Cartan connection $\varpi$ with this $G/H$ Cartan geometry and this mutation can be decomposed as:
where $\varpi_{\mathfrak{h}}^{ab}=\varpi^{ab}_{\mathfrak{h}\>\,\mu}dx^{\mu}=A^{ab}_{~{}\,\,\mu}dx^{\mu}$ corresponds to the spin connection and $\ell\varpi_{\mathfrak{p}}^{a}=\beta^{\,a}_{~{}\,\,\mu}dx^{\mu}$ corresponds to the tetrad while $\varpi_{\mathfrak{p}}=\dfrac{1}{\ell}\beta$ is the soldering form and $\bar{\beta}=\beta^{\text{T}}\eta$. The curvature associated to the covariant derivative and the Cartan connection is:
@@
$\qquad\oplus\dfrac{1}{\ell}\bigl{(}\partial_{\mu}\beta^{\,r}_{~{}\,\,\nu}+\dfrac{1}{4}A^{ab}_{~{}\,\,\mu}\beta^{\,c}_{~{}\,\,\nu}(\eta_{bc}\delta^{r}_{a}-\eta_{ac}\delta^{r}_{b})\bigr{)}dx^{\mu}\wedge dx^{\nu}\otimes M_{r}$ (30)
@@
$T$ $=\ell\bar{\Omega}_{\mathfrak{m}}=d\beta+A\wedge\beta=\bigl{(}\partial_{\mu}\beta^{\,r}_{~{}\,\,\nu}+\dfrac{1}{4}A^{ab}_{~{}\,\,\mu}\beta^{\,c}_{~{}\,\,\nu}(\eta_{bc}\delta^{r}_{a}-\eta_{ac}\delta^{r}_{b})\bigr{)}dx^{\mu}\wedge dx^{\nu}\otimes M_{r}$ (32)
with $\xi=\dfrac{1}{2}\xi^{ab}J_{ab}=\dfrac{1}{4}\xi^{ab}_{~{}~{}\mu\nu}dx^{\mu}\wedge dx^{\nu}\otimes J_{ab}$, where $\xi^{ab}=2\varpi_{\mathfrak{p}}^{a}\wedge\varpi_{\mathfrak{p}}^{b}$ and $\xi^{ab}_{~{}~{}\mu\nu}=2(\varpi^{~{}\,a}_{\mathfrak{p}\,\,\mu}\varpi^{~{}\,b}_{\mathfrak{p}\,\,\nu}-\varpi^{~{}\,a}_{\mathfrak{p}\,\,\nu}\varpi^{~{}\,b}_{\mathfrak{p}\,\,\mu})$. The Bianchi’s identities $D\bar{\Omega}=d\bar{\Omega}+[\varpi,\bar{\Omega}]=0$ split as:
@@ 2.2 Hodge operator
The Hodge operator defined on differential forms on the manifold $\mathscr{M}$ requires a metric $g$ on $\mathscr{M}$. According to the reductive Cartan geometry $G/H$ at hand, $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{m}$, the Killing form $K$ on $\mathfrak{g}$ splits into two parts:
@@
In this case we choose $g=\varpi_{\mathfrak{m}}^{*}K_{\mathfrak{m}}$. It corresponds to the pullback of the $\mathfrak{m}$ part of the Killing form on $\mathfrak{g}$ by $\varpi_{\mathfrak{m}}$ the $\mathfrak{m}$-part of the Cartan connection. For a manifold $\mathscr{M}$ of dimension $m=4$ with Lorentzian Cartan geometry defined by mutation (5):
$K_{\mathfrak{g}}(\varphi,\varphi^{\prime})$ $=K_{\mathfrak{h}}(\varphi_{h},\varphi^{\prime}_{h})+K_{\mathfrak{m}}(\varphi_{m},\varphi^{\prime}_{m})=3\operatorname{Tr}(\varphi_{h},\varphi^{\prime}_{h})+6k\eta(\varphi_{m},\varphi^{\prime}_{m})$
$=-3\eta_{ac}\eta_{bd}\varphi_{\mathfrak{h}}^{ab}\varphi_{h}^{\prime cd}+6k\eta_{ab}\varphi_{\mathfrak{m}}^{a}\varphi_{\mathfrak{m}}^{\prime b}$
@@
$g(X,Y)=(\varpi_{\mathfrak{m}}^{*}K_{\mathfrak{m}})(X,Y)=\dfrac{6k}{\ell^{2}}\eta\bigl{(}\beta(X),\beta(Y)\bigr{)}=\dfrac{6k}{\ell^{2}}\tilde{g}(X,Y)$
where $\tilde{g}=\beta^{*}\eta$ corresponds to the metric usually defined in the tetrad formalism. Let $\omega\in\Omega^{q}(\mathscr{M},\mathfrak{g})$ be a q-form on $\mathscr{M}$ with values in $\mathfrak{g}$. Its local trivialisation in a given chart is :
$\omega=\dfrac{1}{r!}\omega_{\mu_{1}\mu_{2}...\mu_{r}}dx^{\mu_{1}}\wedge dx^{\mu_{2}}\wedge...\wedge dx^{\mu_{r}}$ (2)
Let $\varepsilon_{a_{1}a_{2}...a_{n}}$ be the Levi-Civita symbol of dimension n, such that $\varepsilon_{12...n}=1$. From the metric $g$ one can define a Hodge star operator *** that acts on $\omega$ as :
*$*\omega=$ $\dfrac{1}{r!}\sqrt{|det(g)|}\,\omega_{\mu_{1}...\mu_{r}}g^{\mu_{1}\nu_{1}}...g^{\mu_{r}\nu_{r}}\varepsilon_{\nu_{1}...\nu_{m}}dx^{\nu_{r+1}}\wedge...dx^{\nu_{m}}$ (3)
@@
$dvol=\sqrt{|\tilde{g}|}d^{m}x=\sqrt{|\tilde{g}|}\dfrac{1}{m!}\varepsilon_{\mu_{1}\mu_{2}...\mu_{m}}dx^{\mu_{1}}\wedge dx^{\mu_{2}}\wedge...\wedge dx^{\mu_{m}}=\dfrac{1}{m!}\varepsilon_{a_{1}a_{2}...a_{m}}\beta^{\,a_{1}}\wedge\beta^{\,a_{2}}\wedge...\wedge\beta^{\,a_{m}}.$
@@ 2.3 Associated action
$\varepsilon_{abcd}=\varepsilon_{\mu_{1}\mu_{2}\mu_{3}\mu_{4}}\beta^{-1\,\mu_{1}}_{~{}~{}\,\,\;\;\;a}\beta^{-1\,\mu_{2}}_{~{}~{}\,\,\;\;\;b}\beta^{-1\,\mu_{3}}_{~{}~{}\,\,\;\;\;c}\beta^{-1\,\mu_{4}}_{~{}~{}\,\,\;\;\;d}$ and $\varepsilon^{abcd}=\varepsilon^{\mu_{1}\mu_{2}\mu_{3}\mu_{4}}\beta^{\,a}_{~{}\,\,\mu_{1}}\beta^{\,b}_{~{}\,\,\mu_{2}}\beta^{\,c}_{~{}\,\,\mu_{3}}\beta^{\,d}_{~{}\,\,\mu_{4}}$. The trace $\operatorname{Tr}$ will be used to compute the Pontrjagin number associated to the curvature $\bar{\Omega}$ having values in $\mathfrak{g}$. Here, $\forall\varphi,\varphi^{\prime}\in\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{m}$, $\operatorname{Tr}(\varphi\varphi^{\prime})=-\eta_{ac}\eta_{bd}\varphi_{\mathfrak{h}}^{ab}\varphi_{\mathfrak{h}}^{{}^{\prime}cd}+2k\eta_{ab}\varphi_{\mathfrak{m}}^{a}\varphi_{\mathfrak{m}}^{{}^{\prime}b}$. Taking a linear combination of the Euler and Pontrjagin numbers ($\mathdutchcal{E}$ (8) and $\mathdutchcal{P}$ (10)) associated respectively to $\bar{\Omega}_{\mathfrak{h}}$ and $\bar{\Omega}$ we can build the action :
$S_{G}[\varpi]$ $=e\mathdutchcal{E}+r\mathdutchcal{P}=\int_{\mathscr{M}}\Bigl{(}eE(\mathscr{M})+rP(\bar{\Omega})\Bigr{)}=\int_{\mathscr{M}}\Bigl{(}ePf(\dfrac{\bar{\Omega}_{\mathfrak{h}}}{2\pi})+rP(\bar{\Omega})\Bigr{)}$
$=\int_{\mathscr{M}}\Bigl{(}\dfrac{e}{2(4\pi)^{2}}\varepsilon_{abcd}\bar{\Omega}_{\mathfrak{h}}^{ab}\wedge\bar{\Omega}_{\mathfrak{h}}^{cd}-\dfrac{r}{8\pi^{2}}\operatorname{Tr}(\bar{\Omega}\wedge\bar{\Omega})\Bigr{)}$
@@
$\qquad\quad+\dfrac{r}{8\pi^{2}}\bigl{(}R^{ab}\wedge R_{ab}+2kR^{ab}\wedge\xi_{ab}-\dfrac{2k}{\ell^{2}}T^{a}\wedge T_{a}\bigr{)}\Bigr{)}$
@@
$\qquad\quad+\dfrac{r}{8\pi^{2}}\bigl{(}R^{ab}\wedge R_{ab}+\dfrac{2k}{\ell^{2}}\sqrt{|\tilde{g}|}R^{ab}_{\;\;\;\mu\nu}\beta^{-1\,\mu}_{~{}~{}\,\,\;\;\;c}\beta^{-1\,\nu}_{~{}~{}\,\,\;\;\;d}\varepsilon^{cd}_{~{}~{}ab}d^{4}x-\dfrac{2k}{\ell^{2}}T^{a}\wedge T_{a}\bigr{)}\Bigr{)}.$ (1)
@@
$S_{M}[A,\beta]$ $=\dfrac{\ell^{3}}{216k^{2}}\int_{\mathscr{M}}h^{\varepsilon}\bigl{(}i\psi,K_{\mathfrak{m},ab}\gamma^{a}\varpi_{\mathfrak{m}}^{b}\wedge*D_{\mathfrak{h}}\psi\bigl{)}=ReS_{D}$ (2)
where $ReS_{D}$ corresponds to the real part of the Dirac action as described in [8, p.197]. We are now in position to identify the different terms of $S_{G}[A,\beta]$ above with the action presented in [14] which in our notations correspond to:
@@
where the $\star$ is an internal Hodge such that $\star\beta^{ab}=\dfrac{1}{2}\varepsilon^{ab}_{~{}~{}kl}\beta^{kl}$ as in [12, p.2]. For $G=ISO(3,1)$ ($k=0$) the action simplifies to the topological terms:
$S_{G}[\varpi]$ $=\int_{\mathscr{M}}\Bigl{(}\overbrace{\dfrac{e}{2(4\pi)^{2}}\varepsilon_{abcd}R^{ab}\wedge R^{cd}}^{Euler}+\overbrace{\dfrac{r}{8\pi^{2}}R^{ab}\wedge R_{ab}}^{Pontrjagin}\Bigr{)}$ (4)
In this case from the first Bianchi identity $D\bar{\Omega}_{\mathfrak{h}}=0$, and by neglecting boundary terms, the equations of motion for $A$ and $\beta$ require respectively a null spin density and a null energy-momentum tensor and thus turns out to be physically irrelevant. While for general mutation $\mu$ with parameter $k$, (with particular cases $G^{\prime}=SO(4,1),SO(3,2)$ for $k=-\epsilon=-1,+1$, respectively) we compare $S^{\prime}$ (3) and $S_{G}$ (2.3) and identify the 6 coupling constants present in $S^{\prime}$ via the system:
$\begin{cases}\dfrac{ek}{4\pi^{2}\ell^{2}}=\alpha_{1}\\ \dfrac{rk}{2\pi^{2}\ell^{2}}=(\alpha_{2}-\alpha_{5})\\ \dfrac{r}{8\pi^{2}}=\alpha_{3}\\ \dfrac{e}{(4\pi)^{2}}=\alpha_{4}\\ -\dfrac{rk}{4\pi^{2}\ell^{2}}=\alpha_{5}\\ \dfrac{ek^{2}}{8\pi^{2}\ell^{4}}=\alpha_{6}\end{cases}\Leftrightarrow\begin{cases}\alpha_{1}=\dfrac{4k}{\ell^{2}}\alpha_{4}=\dfrac{2\ell^{2}}{k}\alpha_{6}\\ \alpha_{2}-\alpha_{5}=\dfrac{4k}{\ell^{2}}\alpha_{3}=-2\alpha_{5}\\ \alpha_{1}=\dfrac{2ek}{r\ell^{2}}\alpha_{3}\\ r=\dfrac{2\pi^{2}\ell^{2}}{k}(\alpha_{2}-\alpha_{5})\end{cases}$ (5)
@@
$\begin{cases}\alpha_{1}=\dfrac{4k}{\ell^{2}}\alpha_{4}=\dfrac{2\ell^{2}}{k}\alpha_{6}\\ \alpha_{2}=\dfrac{2k}{\ell^{2}}\alpha_{3}=-\alpha_{5}\\ \alpha_{1}=\dfrac{e}{r}\alpha_{2}\\ r=\dfrac{4\pi^{2}\ell^{2}}{k}\alpha_{2}\end{cases}$ (6)
@@
$S_{G}[\varpi]$ $=e\mathdutchcal{E}+r\mathdutchcal{P}$ (7)
$=\int_{\mathscr{M}}\Bigl{(}\overbrace{\dfrac{k}{4\pi^{2}\ell^{2}}\big{(}\dfrac{e}{2}R^{ab}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}+rR^{ab}\wedge\beta_{a}\wedge\beta_{b}\big{)}}^{Holst}+\overbrace{\dfrac{2ek^{2}}{(4\pi)^{2}\ell^{4}}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}}^{Cosmological~{}constant}$
$\qquad\quad+\overbrace{\dfrac{r}{8\pi^{2}}R^{ab}\wedge R_{ab}}^{Pontrjagin}+\overbrace{\dfrac{e}{2(4\pi)^{2}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}\overbrace{-\dfrac{rk}{4\pi^{2}\ell^{2}}(T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b})}^{Nieh-Yan}\Bigr{)}.$ (8)
@@
$\dfrac{\delta\mathscr{L}_{G}[A,\beta]}{\delta\beta^{c}}$ $=\dfrac{ek}{4\pi^{2}\ell^{2}}(R^{ab}+\dfrac{2k}{\ell^{2}}\beta^{a}\wedge\beta^{b})\wedge\beta^{d}\varepsilon_{abcd}+\dfrac{rk}{\pi^{2}\ell^{2}}R^{ab}\wedge\beta_{b}\delta_{ac}=\tau_{c}=-\dfrac{\delta\mathscr{L}_{M}[A,\beta]}{\delta\beta^{c}}$ (9)
$\dfrac{\delta\mathscr{L}_{G}[A,\beta]}{\delta A^{ab}}$ $=\dfrac{ek}{4\pi^{2}\ell^{2}}T^{c}\wedge\beta^{d}\varepsilon_{abcd}+\dfrac{rk}{\pi^{2}\ell^{2}}T_{a}\wedge\beta_{b}=\dfrac{1}{2}\mathfrak{s}_{ab}=-\dfrac{\delta\mathscr{L}_{M}[A,\beta]}{\delta A^{ab}}$ (10)
where $\tau$ and $\mathfrak{s}$ are respectively the energy-momentum tensor and the spin density of matter. For $r=0$ (no Pontrjagin term), setting $e=\dfrac{8\pi^{2}\alpha_{1}\ell^{2}}{k}=\dfrac{3\pi}{2\Lambda_{0}G}$ and $\ell^{2}=-\dfrac{6k}{\Lambda_{0}}$ where $\alpha_{1}=-\dfrac{1}{16\pi G}$ and applying the Hodge star *** to the first equation (9) gives us the Einstein equations with bare cosmological constant $\Lambda_{0}$ while the other (10) reduces to the equation relating spin-density to torsion of Einstein-Cartan gravity:

### v3

@@ A Cartan connection $\varpi$ with this $G/H$ Cartan geometry and this mutation can be decomposed as:
where $\varpi_{\mathfrak{h}}^{ab}=\varpi^{ab}_{\mathfrak{h}\>\,\mu}dx^{\mu}=A^{ab}_{~{}\,\,\mu}dx^{\mu}$ corresponds to the spin connection and $\ell\varpi_{\mathfrak{p}}^{a}=\beta^{\,a}_{~{}\,\,\mu}dx^{\mu}$ corresponds to the tetrad while $\varpi_{\mathfrak{p}}=\dfrac{1}{\ell}\beta$ is the soldering form and $\bar{\beta}=\beta^{\text{T}}\eta$.
The curvature associated to the covariant derivative and the Cartan connection is:
@@
$\qquad\oplus\dfrac{1}{\ell}\bigl{(}\partial_{\mu}\beta^{\,r}_{~{}\,\,\nu}+\dfrac{1}{2}A^{ab}_{~{}\,\,\mu}\beta^{\,c}_{~{}\,\,\nu}(\eta_{bc}\delta^{r}_{a}-\eta_{ac}\delta^{r}_{b})\bigr{)}dx^{\mu}\wedge dx^{\nu}\otimes M_{r}$ (12)
@@
$T$ $=\ell\bar{\Omega}_{\mathfrak{m}}=d\beta+A\wedge\beta=\bigl{(}\partial_{\mu}\beta^{\,r}_{~{}\,\,\nu}+\dfrac{1}{2}A^{ab}_{~{}\,\,\mu}\beta^{\,c}_{~{}\,\,\nu}(\eta_{bc}\delta^{r}_{a}-\eta_{ac}\delta^{r}_{b})\bigr{)}dx^{\mu}\wedge dx^{\nu}\otimes M_{r}$ (14)
with $\xi=\dfrac{1}{2}\xi^{ab}J_{ab}=\dfrac{1}{4}\xi^{ab}_{~{}~{}\mu\nu}dx^{\mu}\wedge dx^{\nu}\otimes J_{ab}$, where $\xi^{ab}=2\varpi_{\mathfrak{p}}^{a}\wedge\varpi_{\mathfrak{p}}^{b}$ and $\xi^{ab}_{~{}~{}\mu\nu}=2(\varpi^{~{}\,a}_{\mathfrak{p}\,\,\mu}\varpi^{~{}\,b}_{\mathfrak{p}\,\,\nu}-\varpi^{~{}\,a}_{\mathfrak{p}\,\,\nu}\varpi^{~{}\,b}_{\mathfrak{p}\,\,\mu})$ an additional term due to the symmetric algebra structure.
The Bianchi identities $D\bar{\Omega}=d\bar{\Omega}+[\varpi,\bar{\Omega}]=0$ split as:
@@ 2.2 Hodge operator
The Hodge operator defined on differential forms on the manifold $\mathscr{M}$ requires a metric $g$ on $\mathscr{M}$. According to the reductive Cartan geometry $G/H$ at hand, $\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{m}$, [^6] In this Section, one denotes $\mathfrak{g}$ the mutated algebra $\mathfrak{g}^{\prime}$ defined in (1) and $\mathfrak{p}\simeq\mathfrak{m}$. the Killing form $K$ on $\mathfrak{g}$ splits into two parts:
@@
Let $\kappa$ and $\zeta$ be real scalars. In this case we choose $h=\kappa K$ as the metric on the Lie algebra $\mathfrak{g}$ and $g=\zeta\varpi_{\mathfrak{m}}^{*}h=\kappa\zeta\varpi_{\mathfrak{m}}^{*}K_{\mathfrak{m}}$. It corresponds to the pullback of the $\mathfrak{m}$ part of the Killing form on $\mathfrak{g}$ by $\varpi_{\mathfrak{m}}$ the $\mathfrak{m}$-part of the Cartan connection up to the factor $\kappa\zeta$.
For a manifold $\mathscr{M}$ of dimension $m=4$ with Lorentzian Cartan geometry defined by the mutation (1):
$K_{\mathfrak{g}}(\varphi,\varphi^{\prime})$ $=K_{\mathfrak{h}}(\varphi_{\mathfrak{h}},\varphi^{\prime}_{\mathfrak{h}})+K_{\mathfrak{m}}(\varphi_{\mathfrak{m}},\varphi^{\prime}_{\mathfrak{m}})=3\operatorname{Tr}(\varphi_{\mathfrak{h}},\varphi^{\prime}_{\mathfrak{h}})+6k\eta(\varphi_{\mathfrak{m}},\varphi^{\prime}_{\mathfrak{m}})$
$=-3\eta_{ac}\eta_{bd}\varphi_{\mathfrak{h}}^{ab}\varphi_{\mathfrak{h}}^{\prime cd}+6k\eta_{ab}\varphi_{\mathfrak{m}}^{a}\varphi_{\mathfrak{m}}^{\prime b}$
@@
$g(X,Y)=\kappa\zeta(\varpi_{\mathfrak{m}}^{*}K_{\mathfrak{m}})(X,Y)=\dfrac{6k\kappa\zeta}{\ell^{2}}\eta\bigl{(}\beta(X),\beta(Y)\bigr{)}=\dfrac{6k\kappa\zeta}{\ell^{2}}\tilde{g}(X,Y)$
where $\tilde{g}=\beta^{*}\eta$ corresponds to the metric usually defined in the tetrad formalism.
Let $\omega\in\Omega^{r}(\mathscr{M},\mathfrak{g})$ be an $r$-form on $\mathscr{M}$ with values in $\mathfrak{g}$. Its local trivialisation in a given chart is :
$\omega=\dfrac{1}{r!}\omega_{\mu_{1}\mu_{2}...\mu_{r}}dx^{\mu_{1}}\wedge dx^{\mu_{2}}\wedge...\wedge dx^{\mu_{r}}.$ (2)
Let $\varepsilon_{a_{1}a_{2}...a_{n}}$ be the Levi-Civita symbol in dimension $n$, such that $\varepsilon_{12...n}=1$. From the metric $g$ one can define a Hodge star operator * that acts on $\omega$ as :
$*\omega=$ $\dfrac{1}{r!}\sqrt{|det(g)|}\,\omega_{\mu_{1}...\mu_{r}}g^{\mu_{1}\nu_{1}}...g^{\mu_{r}\nu_{r}}\varepsilon_{\nu_{1}...\nu_{m}}dx^{\nu_{r+1}}\wedge...dx^{\nu_{m}}$ (3)
@@
$dvol=\sqrt{|\tilde{g}|}d^{m}x=\sqrt{|\tilde{g}|}\dfrac{1}{m!}\varepsilon_{\mu_{1}\mu_{2}...\mu_{m}}dx^{\mu_{1}}\wedge dx^{\mu_{2}}\wedge...\wedge dx^{\mu_{m}}=\dfrac{1}{m!}\varepsilon_{a_{1}a_{2}...a_{m}}\beta^{\,a_{1}}_{~{}\,\,}\wedge\beta^{\,a_{2}}_{~{}\,\,}\wedge...\wedge\beta^{\,a_{m}}_{~{}\,\,}.$
@@ 2.3 Associated action
The trace $\operatorname{Tr}$ will be used to compute the Pontrjagin number associated to the $\mathfrak{g}$-valued curvature $\bar{\Omega}$. Here, $\forall\varphi,\varphi^{\prime}\in\mathfrak{g}=\mathfrak{h}\oplus\mathfrak{m}$, $\operatorname{Tr}(\varphi\varphi^{\prime})=-\eta_{ac}\eta_{bd}\varphi_{\mathfrak{h}}^{ab}\varphi_{\mathfrak{h}}^{{}^{\prime}cd}+2k\eta_{ab}\varphi_{\mathfrak{m}}^{a}\varphi_{\mathfrak{m}}^{{}^{\prime}b}$.
By combining linearly the Euler number ($\mathdutchcal{E}$ (5)) with the two Pontrjagin numbers ($\mathdutchcal{P}$ (7)) associated to $\bar{\Omega}_{\mathfrak{h}}$ and $\bar{\Omega}$, one can build the action (for $e,r,y$ real numbers):
$S_{G}[\varpi]$ $=e\mathdutchcal{E}(\mathscr{M})+r\mathdutchcal{P}(\bar{\Omega})+y\mathdutchcal{P}(\bar{\Omega}_{\mathfrak{h}})=\int_{\mathscr{M}}\Bigl{(}eE(\mathscr{M})+rP(\bar{\Omega})+yP(\bar{\Omega}_{\mathfrak{h}})\Bigr{)}$
$=\int_{\mathscr{M}}\Bigl{(}e\operatorname{Pf}\left(\tfrac{\bar{\Omega}_{\mathfrak{h}}}{2\pi}\right)+rP(\bar{\Omega})+yP(\bar{\Omega}_{\mathfrak{h}})\Bigr{)}$
$=\int_{\mathscr{M}}\Bigl{(}\dfrac{e}{2(4\pi)^{2}}\varepsilon_{abcd}\bar{\Omega}_{\mathfrak{h}}^{ab}\wedge\bar{\Omega}_{\mathfrak{h}}^{cd}-\dfrac{r}{8\pi^{2}}\operatorname{Tr}(\bar{\Omega}\wedge\bar{\Omega})-\dfrac{y}{8\pi^{2}}\operatorname{Tr}(\bar{\Omega}_{\mathfrak{h}}\wedge\bar{\Omega}_{\mathfrak{h}})\Bigr{)}$
@@
$\qquad\quad+\dfrac{1}{8\pi^{2}}\bigl{(}(r+y)(R^{ab}\wedge R_{ab}+2kR^{ab}\wedge\xi_{ab})-\dfrac{2rk}{\ell^{2}}T^{a}\wedge T_{a}\bigr{)}\Bigr{)}$
@@
$\qquad\quad+\dfrac{1}{8\pi^{2}}\bigl{(}(r+y)(R^{ab}\wedge R_{ab}+\dfrac{2k}{\ell^{2}}\sqrt{|\tilde{g}|}R^{ab}_{\;\;\;\mu\nu}\beta^{-1\,\mu}_{~{}~{}\,\,\;\;\;c}\beta^{-1\,\nu}_{~{}~{}\,\,\;\;\;d}\varepsilon^{cd}_{~{}~{}ab}d^{4}x)-\dfrac{2rk}{\ell^{2}}T^{a}\wedge T_{a}\bigr{)}\Bigr{)}$ (1)
@@
As mentionned in [7, p.192] there exists an isomorphism between elements of the Lorentz Lie algebra $\mathfrak{so}(n-1,1)$ and $\mathfrak{spin}(n-1,1)$ that can be written as $\Phi_{\varepsilon}:\mathfrak{so}(n-1,1)\rightarrow\mathfrak{spin}(n-1,1)$. Its action on the Lie algebra generators of $\mathfrak{so}(n-1,1)$ is:
$\Phi_{\varepsilon}(J_{rs})=\dfrac{1}{4}\bar{e}_{a}J_{rs}e_{b}\gamma^{a}\gamma^{b}=\dfrac{1}{4}(J_{rs})_{ab}\gamma^{a}\gamma^{b}=\dfrac{1}{2}\eta_{ar}\eta_{bs}\gamma^{a}\gamma^{b}.$ (2)
From the covariant derivative on the space of spinors $D_{\mathfrak{h}}=d+\Phi_{\varepsilon}(\varpi_{\mathfrak{h}})=(\partial_{\mu}+\dfrac{1}{4}\varpi_{\mathfrak{h}\>\,ab,\mu}\gamma^{a}\gamma^{b})dx^{\mu}$ associated to $\varpi_{\mathfrak{h}}$ and $h^{\varepsilon}$ one can construct the following matter action: [^7] Since the metrics at the level of the Lie algebra and on $\mathscr{M}$ are always defined up to the factors $\kappa$ and $\zeta$, we set $\kappa^{2}\zeta=\tfrac{\ell^{3}}{216k^{2}}$ such that the overall factor is arranged to stick with the Dirac action given in [7, p.197].
$S_{M}[A,\beta]$ $=\int_{\mathscr{M}}h^{\varepsilon}\bigl{(}i\psi,h_{ab}\gamma^{a}\varpi_{\mathfrak{m}}^{b}\wedge*D_{\mathfrak{h}}\psi\bigl{)}=\mathrm{Re}S_{D}$ (3)
where $\mathrm{Re}S_{D}$ corresponds to the real part of the Dirac action as described in [7, p.197].
We are now in position to identify the different terms of $S_{G}[A,\beta]$ above with the action presented in [8] which, in our notation, reads:
@@
where the $\star$ is an internal Hodge star such that $\star\beta^{ab}=\dfrac{1}{2}\varepsilon^{ab}_{~{}~{}kl}\beta^{kl}$ as in [3, p.2].
For $G=ISO(3,1)$ ($k=0$) the action simplifies to the topological terms:
$S_{G}[\varpi]$ $=\int_{\mathscr{M}}\Bigl{(}\overbrace{\dfrac{e}{2(4\pi)^{2}}\varepsilon_{abcd}R^{ab}\wedge R^{cd}}^{Euler}+\overbrace{\dfrac{r+y}{8\pi^{2}}R^{ab}\wedge R_{ab}}^{Pontrjagin}\Bigr{)}.$ (5)
In this case from the first Bianchi identity $D\bar{\Omega}_{\mathfrak{h}}=0$, and by neglecting boundary terms, the equations of motion for $A$ and $\beta$ require respectively a null spin density and a null energy-momentum tensor and thus turns out to be physically irrelevant.
While for general mutation $\mu$ with parameter $k$, (recall the particular cases $G^{\prime}=SO(4,1),SO(3,2)$ for $k=-\epsilon=-1,+1$, respectively) we compare $S^{\prime}$ (4) and $S_{G}$ (2.3) and identify the 6 coupling constants present in $S^{\prime}$ via the system:
$\begin{cases}\dfrac{ek}{4\pi^{2}\ell^{2}}=\alpha_{1}\\ \dfrac{(r+y)k}{2\pi^{2}\ell^{2}}=(\alpha_{2}-\alpha_{5})\\ \dfrac{r+y}{8\pi^{2}}=\alpha_{3}\\ \dfrac{e}{(4\pi)^{2}}=\alpha_{4}\\ -\dfrac{rk}{4\pi^{2}\ell^{2}}=\alpha_{5}\\ \dfrac{ek^{2}}{8\pi^{2}\ell^{4}}=\alpha_{6}\end{cases}\Leftrightarrow\begin{cases}\alpha_{1}=\dfrac{4k}{\ell^{2}}\alpha_{4}=\dfrac{2\ell^{2}}{k}\alpha_{6}\\ \alpha_{2}-\alpha_{5}=\dfrac{4k}{\ell^{2}}\alpha_{3}=-2(1+\dfrac{y}{r})\alpha_{5}\\ \alpha_{1}=\dfrac{2ek}{(r+y)\ell^{2}}\alpha_{3}\\ r+y=\dfrac{2\pi^{2}\ell^{2}}{k}(\alpha_{2}-\alpha_{5})\end{cases}$ (6)
@@
$\begin{cases}\alpha_{1}=\dfrac{4k}{\ell^{2}}\alpha_{4}=\dfrac{2\ell^{2}}{k}\alpha_{6}\\ \alpha_{2}=\dfrac{k(r+2y)}{4\pi^{2}\ell^{2}}=\dfrac{2k(r+2y)}{(r+y)\ell^{2}}\alpha_{3}=-(1+\dfrac{2y}{r})\alpha_{5}\\ \alpha_{1}=\dfrac{e}{r+2y}\alpha_{2}\end{cases}$ (7)
@@
$S_{G}[\varpi]$ $=e\mathdutchcal{E}(\mathscr{M})+r\mathdutchcal{P}(\bar{\Omega})+y\mathdutchcal{P}(\bar{\Omega}_{\mathfrak{h}})$ (8)
$=\int_{\mathscr{M}}\Bigl{(}\overbrace{\dfrac{k}{4\pi^{2}\ell^{2}}\big{(}\dfrac{e}{2}R^{ab}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}+(r+2y)R^{ab}\wedge\beta_{a}\wedge\beta_{b}\big{)}}^{Holst}+\overbrace{\dfrac{ek^{2}}{8\pi^{2}\ell^{4}}\beta^{a}\wedge\beta^{b}\wedge\beta^{c}\wedge\beta^{d}\varepsilon_{abcd}}^{Cosmological~{}constant}$
$\qquad\quad+\overbrace{\dfrac{r+y}{8\pi^{2}}R^{ab}\wedge R_{ab}}^{Pontrjagin}+\overbrace{\dfrac{e}{2(4\pi)^{2}}R^{ab}\wedge R^{cd}\varepsilon_{abcd}}^{Euler}\overbrace{-\dfrac{rk}{4\pi^{2}\ell^{2}}(T^{a}\wedge T_{a}-R^{ab}\wedge\beta_{a}\wedge\beta_{b})}^{Nieh-Yan}\Bigr{)}.$ (9)
@@
$\dfrac{\delta\mathscr{L}_{G}[A,\beta]}{\delta\beta^{c}}$ $=\dfrac{ek}{4\pi^{2}\ell^{2}}(R^{ab}+\dfrac{2k}{\ell^{2}}\beta^{a}\wedge\beta^{b})\wedge\beta^{d}\varepsilon_{abcd}+\dfrac{(r+y)k}{\pi^{2}\ell^{2}}R^{ab}\wedge\beta_{b}\delta_{ac}=\tau_{c}=-\dfrac{\delta\mathscr{L}_{M}[A,\beta]}{\delta\beta^{c}}$ (10)
$\dfrac{\delta\mathscr{L}_{G}[A,\beta]}{\delta A^{ab}}$ $=\dfrac{ek}{4\pi^{2}\ell^{2}}T^{c}\wedge\beta^{d}\varepsilon_{abcd}+\dfrac{(\tfrac{r}{2}+y)k}{\pi^{2}\ell^{2}}T_{a}\wedge\beta_{b}=\dfrac{1}{2}\mathfrak{s}_{ab}=-\dfrac{\delta\mathscr{L}_{M}[A,\beta]}{\delta A^{ab}}$ (11)
where $\tau$ and $\mathfrak{s}$ are respectively the energy-momentum tensor and the spin density of matter.
For $r=y=0$, setting $e=\dfrac{4\pi^{2}\alpha_{1}\ell^{2}}{k}=\dfrac{3\pi}{2\Lambda_{0}G}$ and $\ell^{2}=-\dfrac{6k}{\Lambda_{0}}$ where $\alpha_{1}=-\dfrac{1}{16\pi G}$ and applying the Hodge star * to the first equation (10) gives us the Einstein equations with bare cosmological constant $\Lambda_{0}$ while the other (11) reduces to the equation relating spin-density to torsion of Einstein-Cartan gravity: